John’s utility function is represented by the following: U(C,L) = (C-400)*(L-100), where C is
expenditure on consumption goods and L is hours of leisure time. Suppose that John receives
$150 per week in investment income regardless of how much he works. He earns a wage of $20
per hour. Assume that John has 110 non-sleeping hours a week that could be devoted to work.
a. Graph John’s budget constraint.
b. Find John’s optimal amount of consumption and leisure.
c. John inherits $300,000 from a relative which gives him $288 per week in additional
investment income. How does this change his hours of work decision? (Calculate C and
L). Is there an income and/or substitution effect? Explain.
d. Now suppose John’s wage increases to $50 per hour but he now still receives only $150
per week in investment income (no inheritance). Find his optimal amount of
consumption and leisure. Is there an income and/or substitution effect? Explain what
we can say about the relative magnitudes, if anything.
Answer #1

UCC, L) (c-400) CL-100 Assumiry mindlaire pe1 Budget constoan C=20 [110-L] +150 teums o [Huu17 C+ 20 L 2350 optfmal bimolle (

Neo Budget constrain 110-L7+150 +288 C= 20 2638 Ct 201 Substitute Pn get we 40L 2638 1600 L =4238105. 15 40 = 20x423 81G00 40

udget constrain C =50 L 110-L]+ 150 50L+ C = 5650 Stopeof -50 Budget Condtraint Neo Optimal conditfon 8 50= C-400 L-10 0 50L-

income motAm 2350+337.5 neuo 2687.5 voge Aate mi- wo20 New constrafnt 2687.8 20L+0 Surstitute in 4OL=2687.S41600 L-4287.S-107